A car dealership has both Electric and Internal Combustion Engine (ICE) vehicles available for sale. 30% of customers looking to purchase a new car are interested in Electric vehicles, while the rest want ICE vehicles. On a particular day, the dealership had 17 people coming in to purchase a new car. What is the probability that the number of people who are interested in an Electric vehicle is more than two standard deviations away from the mean value?


1
Expert's answer
2021-05-24T11:31:34-0400

Let X be the number of people.


"X\sim B(n,p)


\\
X\sim B(17,0.30)\\"


Mean "E(X)=np=17\times 0.3=5.1"


Standard deviation-


"s=\sqrt{np(1-p)}=\sqrt{17\times 0.3\times (1-0.3)}=1.8894"


By using Normal approximation of Binomal distribution.


"X\sim N(\mu=5.1,\sigma=1.8894)"


Probability that the number of people who are interested in an Electric vehicle is more than two standard deviations away from the mean value is-


"=P(X<\mu-2\sigma)+P(X>\mu+2\sigma)

\\[9pt]

 =P(Z<-2)+P(z>2)
\\[9pt]


=0.0228+0.0228\\[9pt]=0.0456"